A swimming pool 100 metres long and 40 metres wide is 1 metre deep at the shallow end and 5 metre at the deep end. Find the volume (in cubic metres) of water contained in the pool.
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

100*40(1+5)/2
4,000*6/2
4,000*3=12,000 CUBIC METERS OF WATER.

Answer by Alan3354(69443)

A swimming pool 100 metres long and 40 metres wide is 1 metre deep at the shallow end and 5 metre at the deep end. Find the volume (in cubic metres) of water contained in the pool.
-------------------
Since you didn't specify otherwise, I'll assume the bottom is flat (not level, but flat)
That makes the average depth 3 meters (1+5)/2
Volume = 100*40*3 = 12000 cubic meters
---------------
That's a large pool, 330 feet long. Have you seen it?